OpenStudy (wibrahim):

help plz boys. Consider the equation 3tan^2(x)-1=0 The solutions of this equation over the interval [0, π] are a) 2pi/3, 5pi/6 b) pi/6, 2pi/3 c) 7pi/6, 11pi/6 d) pi/6, 5pi/6 plz help

OpenStudy (math&ing001):

What do you get when solving for tan(x) ?

OpenStudy (cruffo):

I get \[\frac{\sqrt{3}}{3}\]

OpenStudy (math&ing001):

Should be 1/sqrt(3)

OpenStudy (math&ing001):

or -1/sqrt(3)

OpenStudy (cruffo):

yes, but rationalize the denominator :)

OpenStudy (math&ing001):

Lol I didn't notice :P

OpenStudy (math&ing001):

But you forgot about -sqrt(3)/3

OpenStudy (cruffo):

right .. you got me ..., plus or minus

OpenStudy (math&ing001):

Yeah, after that you just apply the inverse function in a calculator..

OpenStudy (math&ing001):

It's called arctan

OpenStudy (chevyrebel04):

ot (-5pi/12) well that will be in quadrant 4 since it is clockwise from the x axis due to the minus sign and 5 pi/12 is less than pi/2 or 6 pi/12. call the x component of this vector a and the y component b. Then the magnitude of the cotan is a/b. Notice that the angle to the negative y axis is 6pi/12 - 5pi/12 = pi/12 then the magnitude of tan pi/12 is a/b, just what we are looking for so what is tan pi/12? In degrees it is 15 degrees. We know tan 30 degrees = 1/sqrt3 and sin 30 =1/2 and cos 30 = sqrt3 /2 so what is tan 15 degrees? well tan 15 = sin 30/(1+cos 30) I think you can take it from there.

OpenStudy (cruffo):

what... calculator.... just use the unit circle. \[ \frac{\sqrt{3}}{3} = \frac{\frac{1}{2}}{\frac{\sqrt{3}}{2}}\] so when is sin(x) =1/2

OpenStudy (cruffo):

noticed @wibrahim is offline...so I guess this is just us

OpenStudy (math&ing001):

Yeah, we'll continue though for future reference ;]

OpenStudy (cruffo):

I'm getting pi/6 and 5pi/6

OpenStudy (math&ing001):

Yep, correct !